# Simulation of Polynomial-Time Nondeterministic Turing Machines via Universal Nondeterministic Turing Machines

### v17

@@ 1. Introduction
where $\overline{L}$ is the set-theoretic complement of $L$ (i.e., assume $L\subseteq\{0,1\}^{*}$, then $\overline{L}=\{0,1\}^{*}\setminus L$). The outstanding and long-standing open conjecture (i.e., the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem) asks whether $\mathcal{NP}$ equals ${\rm co}\mathcal{NP}$ or not and is a central topic in computational complexity theory, whose implications are profound for computer science, mathematics, and related fields.
@@
So far, although the exact relationship between the complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ is unknown and the classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ are widely believed to be different, Figure 1 below illustrates the possibility of the relationship between $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ that most complexity theorists believed.
@@ 1.1. Our Contributions
Note that we prove the first half of Theorem 1.1 by the ordinary simulation techniques and the relationship between languages $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and the complement $\overline{L}(M)$ of $L(M)$. Furthermore, since, by the result of [BGS75], there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, and just as the result that $\mathcal{P}^{A}=\mathcal{NP}^{A}$ [BGS75] led to a strong belief that problems with contradictory relativization are very hard to solve and are not amenable to current proof techniques, i.e., the solutions of such problems are beyond the current state of mathematics (see e.g. [HCCRR93, Hop84]), the conclusion $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ also suggests that the problem of separating $\mathcal{NP}$ from ${\rm co}\mathcal{NP}$ is beyond the current state of mathematics. So, the next result, i.e., the following theorem, is to break the so-called Relativization Barrier:
@@
Meanwhile, there is another case, i.e., $\mathcal{P}^{A}=\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$. In this scenario, we remark that Theorem 6.1 in Section 6 perfectly expounds the reason why we are unable to apply simulation techniques to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$.
@@ 1.2. Background and Prior Work
Computational complexity theory is a central subfield of theoretical computer science and mathematics, which mainly concerns the efficiency of Turing machines (i.e., algorithms) or the intrinsic complexity of computational tasks, i.e., focuses on classifying computational problems according to their resource usage and relating these classes to each other (see e.g. [A1]). In other words, computational complexity theory specifically deals with fundamental questions such as what is feasible computation and what can and cannot be computed with a reasonable amount of computational resources in terms of time or space (memory). What’s exciting is that there are many remarkable open problems in the area of computational complexity theory, such as the famous $\mathcal{P}$ versus $\mathcal{NP}$ problem and the important unknown relation between the complexity class $\mathcal{NP}$ and the complexity class ${\rm co}\mathcal{NP}$, which deeply attract many researchers to conquer them. However, despite decades of effort, very little progress has been made on these important problems. Furthermore, as pointed out by Wigderson [Wig07], to understand the power and limits of efficient computation has led to the development of computational complexity theory, and this discipline in general, and the aforementioned famous open problems in particular, have gained prominence within the mathematics community in the past decades.
@@ 1.4. Organization
The remainder of this paper is organized as follows: For the convenience of the reader, we will review some notions in the next section closely associated with our discussions and fix some notation we will use in the following context, with some useful technical lemmas also presented in the next section. In Section 3, we prove that all polynomial-time nondeterministic Turing machines are enumerable and all ${\rm co}\mathcal{NP}$-machines are enumerable, thus showing that there exists an enumeration of all ${\rm co}\mathcal{NP}$-machines. Section 4 contains the definition of our universal nondeterministic Turing machine, which accepts a language $L_{s}$ not in ${\rm co}\mathcal{NP}$, showing the desired lower bounds. And the Section 5 is devoted to showing that the language $L_{s}$ is in fact in $\mathcal{NP}$, proving the upper bounds required, in which we finish the proof of Theorem 1.1 at our own convenience. In Section 6 we will show Theorem 1.2 to break the so-called Relativization Barrier. In Section 7, we show that there are ${\rm co}\mathcal{NP}$-intermediate languages in the complexity class ${\rm co}\mathcal{NP}$. In Section 8, we prove the Theorem 1.8 which says that no Frege proof systems are polynomially bounded, thus answering an important open question in the area of proof complexity. Finally, we list some open problem for future research in the last section.
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
One may wish to apply simulation techniques to the case of $\mathcal{P}$ versus ${\rm co}\mathcal{P}$ and doubt that simulation techniques are unable to produce a language not in ${\rm co}\mathcal{P}$. In this scenario, we remark that the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}$ via a universal deterministic Turing machine $U^{\prime}$, because a polynomial-time deterministic Turing machine ($\mathcal{P}$ machine) under the “there exists no accepting path” condition is still a polynomial-time deterministic Turing machine. Namely, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic Turing machines are symmetrical. Specifically, for an arbitrary deterministic Turing machine $M$, if $M$ accepts the input $w$, then there is only one accepting path, and if $M$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic Turing machines. In short, if we apply the same technique, i.e., the simulation technique, to create a language not in ${\rm co}\mathcal{P}$, then the universal deterministic Turing machine $U^{\prime}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{\prime})\in\mathcal{P}$ (the proof is similar to the case in Section 5 below). But by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic Turing machines, we clearly have that $L(U^{\prime})\in{\rm co}\mathcal{P}$. Thus, we are unable to claim that $L(U^{\prime})\notin{\rm co}\mathcal{P}$.
@@ 6. Breaking the Relativization Barrier
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to the Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.

### v18

@@ 1. Introduction
where $\overline{L}$ is the set-theoretic complement of $L$ (i.e., assume $L\subseteq\{0,1\}^{*}$, then $\overline{L}=\{0,1\}^{*}\setminus L$). The outstanding and long-standing open problem $\mathcal{NP}\overset{?}{=}{\rm co}\mathcal{NP}$ asks whether $\mathcal{NP}$ equals ${\rm co}\mathcal{NP}$ or not (i.e., the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem) and is a central topic in computational complexity theory, whose implications are profound for computer science, mathematics, and related fields.
@@
Thus far, although the exact relationship between the complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ is unknown and the complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ are widely believed to be different, Figure 1 below illustrates the possibility of the relationship between the complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ that most complexity theorists believed.
@@ 1.1. Our Contributions
Note that we prove the first half of Theorem 1.1 by the ordinary simulation techniques and the relationship between languages $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and the complement $\overline{L}(M)$ of $L(M)$. Furthermore, by the result of [BGS75], since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, and just as the result that $\mathcal{P}^{A}=\mathcal{NP}^{A}$ [BGS75] led to a strong belief that problems with contradictory relativization are very hard to solve and are not amenable to current proof techniques, i.e., the solutions of such problems are beyond the current state of mathematics (see e.g. [HCCRR93, Hop84]), the conclusion $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ also suggests that the problem of separating $\mathcal{NP}$ from ${\rm co}\mathcal{NP}$ is beyond the current state of mathematics. Therefore, the next result, i.e., the following theorem, is to break the so-called Relativization Barrier:
@@
Meanwhile, there is another case, i.e., the case that $\mathcal{P}^{A}=\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$. In this scenario, we remark that Theorem 6.1 in Section 6 perfectly expounds the reason why we are unable to apply simulation techniques to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$.
@@ 1.2. Background and Prior Work
Why are some problems easy for computers to solve, while others seem fundamentally intractable? This question lies at the heart of computational complexity theory. So what is computational complexity theory? Computational complexity theory is a central subfield of theoretical computer science and mathematics, which mainly concerns the efficiency of Turing machines (i.e., algorithms) or the intrinsic complexity of computational tasks, i.e., focuses on classifying computational problems according to their resource usage and relating these classes to each other (see e.g. [A1]). In other words, computational complexity theory specifically deals with fundamental questions such as what is feasible computation and what can and cannot be computed with a reasonable amount of computational resources in terms of time or space (memory). What’s exciting is that there are many remarkable open problems in the area of computational complexity theory, such as the famous $\mathcal{P}$ versus $\mathcal{NP}$ problem and the important unknown relation between the complexity class $\mathcal{NP}$ and the complexity class ${\rm co}\mathcal{NP}$, which deeply attract many researchers to conquer them. However, despite decades of effort, very little progress has been made on these important problems. Furthermore, as pointed out by Wigderson [Wig07], to understand the power and limits of efficient computation has led to the development of computational complexity theory, and this discipline in general, and the aforementioned famous open problems in particular, have gained prominence within the mathematics community in the past decades.
@@ 1.4. Organization
The remainder of this paper is organized as follows: For the convenience of the reader, we will review some notions in the next section closely associated with our discussions and fix some notation we will use in the following context, with some useful technical lemmas also presented in the next section. In Section 3, we prove that all polynomial-time nondeterministic Turing machines are enumerable and all ${\rm co}\mathcal{NP}$-machines are enumerable, thus showing that there exists an enumeration of all ${\rm co}\mathcal{NP}$-machines. Section 4 contains the definition of our universal nondeterministic Turing machine, which accepts a language $L_{s}$ not in ${\rm co}\mathcal{NP}$, showing the desired lower bounds. And the Section 5 is devoted to showing that the language $L_{s}$ is in fact in $\mathcal{NP}$, proving the upper bounds required, in which we finish the proof of Theorem 1.1 at our own convenience. In Section 6 we will show Theorem 1.2 to break the so-called Relativization Barrier. In Section 7, we show that there are ${\rm co}\mathcal{NP}$-intermediate languages in the complexity class ${\rm co}\mathcal{NP}$. In Section 8, we prove the Theorem 1.8 which says that no Frege proof systems are polynomially bounded, thus answering an important open question in the area of proof complexity. Finally, we list some open problems for future research in the last section.
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
Remark 3.1.
There is another way to enumerate all polynomial-time nondeterministic Turing machines without encoding the order of polynomial into their representation. To do so, we need the Cantor pairing function (see Fig. 2 below from [A8]):
$\pi:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$
defined by
$\pi(k_{1},k_{2}):=\frac{1}{2}(k_{1}+k_{2})(k_{1}+k_{2}+1)+k_{2}$
where $k_{1},k_{2}\in\mathbb{N}$. Since the Cantor pairing function is invertible (see [A8]), it is a bijection between $\mathbb{N}\times\mathbb{N}$ and $\mathbb{N}$. Note that we can further define a bijection $\tau$ between $\mathbb{N}$ and $\mathbb{N}_{1}$ such as $\tau:k\mapsto k+1$ where $k\in\mathbb{N}$ and $k+1\in\mathbb{N}_{1}$, which shows that there exists a $(1,1)$ correspondence between $\mathbb{N}$ and $\mathbb{N}_{1}$. As we have shown that any polynomial-time nondeterministic Turing machine is an integer, we can place any polynomial-time nondeterministic Turing machine and its order of polynomial into a tuple $(M,k)$ and use the Cantor pairing function to map the tuple $(M,k)$ to an integer in $\mathbb{N}_{1}$. The reader can easily check that such a method gives an enumeration of the set of polynomial-time nondeterministic Turing machines.
Figure 2. Cantor pairing function
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
One may wish to apply simulation techniques to the case of $\mathcal{P}$ versus ${\rm co}\mathcal{P}$ and doubt that simulation techniques are unable to produce a language not in ${\rm co}\mathcal{P}$. In this scenario, we remark that the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}$ via a universal deterministic Turing machine $U^{\prime}$, because a polynomial-time deterministic Turing machine ($\mathcal{P}$ machine) that works under the “there exists no accepting path” condition is still a polynomial-time deterministic Turing machine. Namely, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic Turing machines are symmetrical. Specifically, for an arbitrary deterministic Turing machine $M$, if $M$ accepts the input $w$, then there is only one accepting path, and if $M$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic Turing machines. In short, if we apply the same technique, i.e., the simulation technique, to create a language not in ${\rm co}\mathcal{P}$, then the universal deterministic Turing machine $U^{\prime}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{\prime})\in\mathcal{P}$ (the proof is similar to the case in Section 5 below). But by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic Turing machines, we clearly have that $L(U^{\prime})\in{\rm co}\mathcal{P}$. Thus, we are unable to claim that $L(U^{\prime})\notin{\rm co}\mathcal{P}$.
@@ 6. Breaking the Relativization Barrier
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to the Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine that works under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.
@@ 9. Open Problems
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contains the problem Satisfiability? Note that by footnote $1$, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this can not exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. If one can show that Satisfiability is not in $\mathcal{BQP}$, then the relationship between $\mathcal{BQP}$ and $\mathcal{NP}$ were clear. Currently, we have no answer to this problem. We tried to think about it but also failed. We remark that this is also a fascinating open problem in complexity theory.
Appendix A Two Definitions of $\mathcal{NP}$ Are Equivalent
The official descriptions on the definitions of $\mathcal{NP}$ in [Coo00] is given by:
Definition A.1.
$\mathcal{NP}\overset{{\rm def}}{=}\{L\,:\,w\in L\Leftrightarrow\exists y(|y|\leq|w|^{k}\,\,\,\text{and }R(w,y))\}$
where the checking relation
$R(w,y)$
is polynomial-time, i.e., the language
$L_{R}\overset{{\rm def}}{=}\{w\#y\,:\,R(w,y)\}$
is in $\mathcal{P}$.
The complexity class $\mathcal{P}$, which is referred to by Definition A.1, is defined as follows:
Definition A.2.
$\begin{array}[]{ll}\mathcal{P}\overset{{\rm def}}{=}&\{L\,:\,L=L(M)\text{ for some Turing machine $M$ that runs}\\ &\qquad\text{ in polynomial time }\}.\end{array}$
In Definition A.2, Turing machine $M$ refers to a deterministic Turing machine, and the notion of “running in polynomial time” is that if there exists $k$ ($\in\mathbb{N}_{1}$) such that,
$T_{M}(n)\leq n^{k}+k,\quad\forall n\in\mathbb{N}_{1}$
where $T_{M}(n)$ is defined by
$T_{M}(n)=\max\{t_{M}(w)\,:\,w\in\Sigma^{n}\}$
and $t_{M}(w)$ denotes the number of steps (or moves) in the computation of $M$ on input $w$ of length $n$.
As you may note that, this definition (i.e., Definition A.1) looks a little different from that given in Section 2. But they are equivalent. Here, we will show the equivalence for these two definitions of $\mathcal{NP}$.
A.1. A Proof of Two Definitions of $\mathcal{NP}$ Are Equivalent.
The remainder of this Appendix is to show that Definition A.1 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent.
To begin, we should first give another definition of $\mathcal{NP}$ which is similar to the official descriptions of the definition of $\mathcal{P}$ in [Coo00] (i.e., Definition A.2 above):
Definition A.3.
$\begin{array}[]{ll}\mathcal{NP}\overset{{\rm def}}{=}&\{L\,:\,L=L(M)\text{ forsome nondeterministic Turing machine $M$}\\ &\qquad\text{that runs in polynomial time}\}.\end{array}$
Now, let us finish the last step of the task that we mentioned at the beginning of this subsection.
Proof A.4.
We need to show that Definition A.3 is equivalent to Definition A.1 and to show the equivalence between Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2. By a simple argument, it is clear that Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent. We proceed as follows:
We show first the “if” part. Suppose that
$L\in\bigcup_{i\in\mathbb{N}_{1}}{\rm NTIME}[n^{i}].$
Then there is a $k\in\mathbb{N}_{1}$, such that
$L\in{\rm NTIME}[n^{k}]$
which means that for all $n\in\mathbb{N}_{1}$, there is a nondeterministic Turing machine $M$, for any $w\in\Sigma^{n}$
$T_{M}(|w|)\leq c_{0}n^{k}+c_{1}n^{k-1}+\cdots+c_{k-1}n+c_{k}\quad\text{where $c_{0}>0$}.$
and
$L=L(M).$
For any constants $c_{0}$, $c_{1}$, $\cdots$, $c_{k}$, there must exists a minimal $t\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}$ and for any $w\in\Sigma^{n}$
$T_{M}(|w|)\leq n^{t}+t.$
So,
$L\in\{\mathcal{L}\,:\,\mathcal{L}=L(M)\text{ for some nondeterministic Turing machine $M$ that runs in polynomial time}\}.$
We show next the “only if” part. Suppose now that the language
$L\in\{\mathcal{L}\,:\,\mathcal{L}=L(M)\,\text{ for some nondeterministic Turing machine $M$ that runs in polynomial time}\}.$
Then there exists a $k\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}_{1}$ and for all $w\in\Sigma^{n}$,
$T_{M}(|w|)\leq n^{k}+k$
which implies that
$L\in{\rm NTIME}[n^{k}]\subseteq\bigcup_{i\in\mathbb{N}_{1}}{\rm NTIME}[n^{i}].$
It is time for us to show that Definition A.3 is equivalent to Definition A.1. For the purposes of brevity, we omit the proof and refer the reader to [Kar72] (cf. Theorem 1 in [Kar72]), or to the proof of Theorem $7$.20202020 in [Sip13] which is a theorem answering the same question discussed here (see p. $294$ of [Sip13]). Thus this completes the proof.